What is A=a+(n-1)d and it says to solve for n
Answer by AnlytcPhil(1810)

What is A = a + (n - 1)d and it says 
to solve for n

This is a formula for the nth term, A, 
of an arithmetic sequence with first
term a and common difference d.

To solve it for n:

         A = a + (n - 1)d

         A = a + d(n - 1)

         A = a + dn - d

A - a + d = dn

Divide both sides by d

 A - a + d     dn
----------- = ----
     d         d


               1
 A - a + d     dn
----------- = ----
     d         d
               1


 A - a + d     
----------- = n
     d         

               A - a + d     
          n = ----------- 
                   d  

Edwin